Find x by solving for the system of equations using any method you prefer.
-8x + 4y = 9 and 3x = 4y + 6
Please help me !

Found 2 solutions by sachi, lenny460:
Answer by sachi(548)

rewriting both the questions we have
-8x + 4y = 9 ..........1
3x - 4y = 6...........2
adding both eqns
-8x+3x=9+6
or -5x=15
or x=-3
from eqn 2
-3*3-4y=6
or 4y=-9-6=-15
or y=-15/4
ans

Answer by lenny460(1073)

Find x by solving for the system of equations using any method you prefer.
-8x + 4y = 9 and 3x = 4y + 6
Please help me !

-8x + 4y = 9 (Equ 1)
3x = 4y + 6 (Equ 2)

3x = 4y + 6 (Equ 2)
Subtract 4y from each side of the equation.
3x - 4y
4y + 6 - 4y
Combine like terms:
4y - 4y = 0
3x - 4y = 6 (Equ 3)
Add Equ 1 and Equ 3
-8x + 4y + 3x - 4y = 9 + 6
Combine like terms:
-8x + 3x = -5x
4y - 4y = 0
9 + 6 = 15
-5x = 15
Divide each side by -5
-5x/-5 = 15/-5
x = -3

Substitute x = -3 in Equ 3
3x - 4y = 6
3(-3) - 4y = 6
3(-3) = -9
-9 - 4y = 6
Add +9 to each side of the equation.
-9 - 4y + 9
6 + 9 = 15
Combine like terms:
-9 + 9 = 0
-4y = 15
Divide each side by -4
-4y/-4 = 15
y = -15/4

Check:
-8x + 4y = 9
-8(-3) + 4(-15/4) = 9
-8 * -3 = 24
4 * -15/4 = -60/4 = -15
24 - 15 = 9
9 = 9

Therefore;
x -3 and
y = -15/4



Lennox Obuong
Algebra 2 Student
Email: obuong3@aol.com